curlyheadedluis:

A tsunami is a destructive, fast-moving ocean wave that is caused by an undersea earthquake, landslide, or volcano. The Pacific Tsunami Warning Center is responsible for monitoring earthquakes that could potentially cause tsunamis in the Pacific Ocean. Through measuring the water level and calculating the speed of a tsunami, scientists can predict the arrival times of tsunamis. The speed, s (in meters per second) at which a tsunami moves is determined by the depth, d (in meters) of the ocean. s=gd−−√, where g is approximately 14.5 meters per second squared. Find the depth of a tsunami in a region of the ocean that is moving at the average speed of 92 meters per second. Round your answer to the nearest meter. Do not include units in your answer.

curlyheadedluis:

need help

0mega:

alright you're posting something you already answered

curlyheadedluis:

and its wrong

0mega:

how are we supposed to help with something, when it gives the correct answer right there

curlyheadedluis:

i wanna know how to do it

curlyheadedluis:

like steps to get the right anwser next time

ghostmasterTy:

you still need help

curlyheadedluis:

mhm

ghostmasterTy:

yes or no

curlyheadedluis:

yes

ghostmasterTy:

ok the whole thing

ghostmasterTy:

or what part

curlyheadedluis:

yes please i just need steps

curlyheadedluis:

like on how to solve it imma a freshman btw

ghostmasterTy:

ok step one is

ghostmasterTy:

read it all of it did you do that

curlyheadedluis:

yes

curlyheadedluis:

ok what abt tht number

curlyheadedluis:

whats the anwser

curlyheadedluis:

wrong

curlyheadedluis:

it was 584

AZ:

Since you already have the answer and you just want to know how to do this question, I'll just work it out for you since it's already been more than 2 days since someone replied to this post. The formula is \( s = \sqrt{gd}\) s is speed in m/s d is depth in m and g is approximately 14.5 m/s^2

AZ:

They ask us when the speed is 92, what is the depth of the tsunami? So plug in 92 for 's' and 14.5 for 'g' and then solve for 'd' \( s = \sqrt{gd}\) \( 92 = \sqrt{14.5d}\) Square both sides \( 92^2 = (\sqrt{14.5d})^2\) and then we get \(8464 = (\sqrt{14.5d})^2\) \( 8464 = 14.5d\) and then to solve for 'd', you just divide both sides by 14.5 and you get d = 583.7 meters which can be rounded to 584